OpenStudy (anonymous):

Two cars leave town, one driving North the other East. They are 104 miles apart when one of them is 56 miles farther from town than the other. At that time how far were they from each town?

OpenStudy (anonymous):

First car: 104^2 = x^2 + (x+56)^2 x = 40 Second car: 40+56=96

OpenStudy (zepp):

 with pythagore you should be able to solve it... ^